# Game Theory and Multi-Agent Reinforcement Learning for Zonal Ancillary Markets

### v3

@@ III Theoretical Framework
In a large part of the literature, the convexity of the agents’ optimization problems is an essential assumption under which noncooperative games are studied and analyzed. On the contrary, in this paper, $\Gamma$ is a generalized Nash game with side constraints, involving trilinear objective functions and nonconvex constraints thus giving rise to a nonconvex game. In order to proceed we will rely on results established in [18], where the authors study a class of problems with similar settings. To that purpose, we first introduce a penalized version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version; then we present the settings of [18], which we use to prove existence of market equilibrium under mild assumptions. Finally we highlight the potential structure of $\Gamma$.
III-A Penalized Generalized Nash Game
We consider a penalized version of our problem, which makes it more tractable and it will allow us to use mild assumptions to obtain an equilibrium existence guarantee. We add the constraints in Eqs. (3) and (6) to the objective function, which gives the function $J_{n}^{p}(w)$ and the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$ (see Appendix B for more details). We can formulate the penalized generalized Nash game $\Gamma^{p}$ as follows:
$\forall n\in\mathcal{N},\quad$ $\min_{w_{n}\in\mathcal{W}_{n}}$ $-J_{n}^{p}(w)$ (14a)
@@ III-A Penalized Generalized Nash Game
If $K_{n}\geq\frac{Z}{3}$, there exists a set of penalties such that a solution of the penalized game (14) is a solution of $\Gamma$.
Proposition 1 results from the comparisons of the sets of KKTs of the generalized Nash game $\Gamma$ and the penalized game (14) denoted $\Gamma^{p}$. Details can be found in Appendix C.
Thanks to this reformulation we obtain copositive Hessians matrices associated to the constraints $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, which is the key point in obtaining an existence result. We discuss the form of these matrices in Appendix D.
III-B Nonconvex Game with Side Constraints
The formulation of [18] considers a problem where the coupling constraints are treated as penalties with associated prices, in a modified objective function. This problem is then solved as an $(N+1)$-player game, where $N$ players optimize their variables and an external one optimizes the prices, which are shared by all players.
With the compact notation previously introduced, we can write the problem in Eq. (14) as an $(N+1)$-player game with prices [^1] Notice that to directly extend the results of [18], we now work with the negative of the objective function and in minimization settings.:
$\forall n\in\mathcal{N},\quad\min_{w_{n}\in\widetilde{\mathcal{W}}^{p}_{n}}\left[-J_{n}^{p}(w_{n},w_{-n})+\sum_{j=1}^{r}\rho_{j}G_{j}^{p}(w)\right]$ (15)
where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$ and introduce the penalized local feasibility set $\widetilde{\mathcal{W}}_{n}^{p}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}^{p}(w_{n})\leq 0\}$. We can now recall the definition of Nash Equilibrium [18, Definition 1] that is a relevant solution concept to analyze this class of games:
Definition 1 (Nash Equilibrium with Side Constraints).
A Nash equilibrium (NE) for the game (15) is a strategy-price tuple $(w^{\star},\rho^{\star})$, such that for all $n$:
$w_{n}^{\star}\in\operatorname*{argmin}_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[-J_{n}(w_{n},w_{-n}^{\star})+\sum_{j=1}^{r}\rho_{j}^{\star}G_{j}(w_{n},w_{-n}^{\star})\right],$
and
$0\leq\rho^{\star}\perp G(w^{\star})\leq 0.$
@@ III-C Existence of a Generalized Nash Equilibrium for $\Gamma$
To obtain existence result for the penlaized problem, we have to verify assumptions on the variables’ domain, the objective functions and the constraint sets. Thanks to our penalized reformulation, we only need to verify a set of mild assumptions:
@@
    $J_{n}(\cdot,w_{-n})$ is continuously differentiable in $w_{n}$, $h_{n}(\cdot)$ and $G(\cdot,w_{-n})$ are twice continuously differentiable in $w_{n}$ for all $n$;
@@
    There exists $w^{\sharp}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h_{n}(w_{n}^{\sharp})<0,\,\forall n\in\mathcal{N}$ and $G(w^{\sharp})<0$;
Theorem 1 (Existence of a solution for $\Gamma$).
Define $\mathcal{W}^{\star}\subseteq\mathcal{W}$ such that for any $w\in\mathcal{W}^{\star}$, the Hessian matrix of the objective function is positive semi-definite. Then, there exists a generalized Nash equilibrium on $\mathcal{W}^{\star}$ solution of $\Gamma^{p}$.
@@
Assumption 1 (a) holds, since $\mathcal{W}_{n}$ is defined as a polyhedral and compact set, i.e., we can express any condition of the form $x\in[x_{a},x_{b}]$ as a set of inequalities $x\leq x_{b},x\geq x_{a}$, which are defined as intersections between closed half spaces. Further, the objective function is continuous and differentiable everywhere in the domain $\mathcal{W}^{\star}$, and the constraints functions are bilinear or trilinear. Thus, they are twice continuous and differentiable over the whole domain $\mathcal{W}^{\star}$. Assumption 1 (b) is then satisfied. Assumption 1 (c) requires the existence of an interior point. Considering $w^{\sharp}$ that strictly satisfies Eq. (3), (4), (5) and (6), we immediately obtain that $\lambda_{n}^{z}=0,\mu_{n,k}=0,\sigma_{n}=0,\delta_{n}=0,\,\forall k\in[1,K_{n}],\,\forall n\in\mathcal{N},\,\forall z\in\mathcal{Z}$. Eq. (12c) needs instead to be relaxed to a strict inequality. We then apply [18, Theorem 13] to the $(N+1)$-players game to obtain the existence of a generalized Nash equilibrium for $\Gamma^{p}$. ∎
Corollary 1.
The generalized Nash game $\Gamma$ defined in (13) admits a solution.
Proof.
Using results from [18, Sec. 2.1], we have an equivalence between the Nash equilibrium of the $(N+1)$-players game and the associated generalized Nash game, which in our case is the penalized version $\Gamma^{p}$. We conclude using Proposition 1, which states the equivalence between $\Gamma^{p}$ and $\Gamma$. ∎
@@
Using [19, Def. 2.1], two requirements need to be satisfied for a GNG to be a GPG. The first requirement amounts to have a feasible set that is nonempty, which is easily satisfied in our setting. In particular, the feasible set, which we call $C$, does not have to be the cartesian product of the players’ feasible sets, meaning that $C\subseteq\prod_{n=1}^{N}\bar{\mathcal{W}}_{n}(w_{-n})$, where $\bar{\mathcal{W}}_{n}(w_{-n}):=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0,G(w)\leq 0\}$. The second requirement is for a potential $P(w)$ function to exist, which in our case is satisfied by using $P(w):=\sum_{n=1}^{N}J_{n}(w)$. ∎
From Proposition 2, we can rewrite Eqs. (13) as:
@@
$\quad G(w_{n},w_{-n})\leq 0,\,\forall n,$ (16c)
which can be solved using a standard nonlinear programming solver.
@@ IV Algorithmic Solutions
We solve the market problem relying on integrated optimization and Gauss-Seidel best-response method, that we compare against a MARL algorithm. For the integrated optimization approach we use the MINOS solver, which relies on a projected Lagrangian method [20]. Next, we discuss the best-response algorithm we use and how we built the multi-agent simulation.
IV-A Gauss-Seidel Best Response
We follow the implementation of the Gauss-Seidel Best Response algorithm from [19]. Here we remove the coordination between players, and we solve the problem in Eq. (13) for each agent iteratively (once again with the MINOS solver), stopping the whole process once the decision variables of the agents do not change anymore, as summarized in Alg. 1.
@@ IV-A Gauss-Seidel Best Response
1:Initialize problem with feasible point $w^{0}=[w_{1}^{0},\dots,w_{N}^{0}]$, $k=0$
2:while Termination criterion is false do
3:  for $n\in[1,\dots,N]$ do
@@
The integrated solution can be seen as a coordinated solution for all the producers at once, while using the Best-Response algorithm allows for each producer to optimize its strategy alone. In the results section we will explain more in details how these two solutions differ in modeling the interactions between agents.
Furthermore, from [19, Theorem 5.2] and the potential structure of our game, we have a guarantee of convergence to a Nash equilibrium, provided it exists, which we already proved in Section III-C.
These two approaches allow us to compare fully coordinated and uncoordinated interaction schemes between agents, that we will compare against the multi-agent simulation based on reinforcement learning.
IV-B Multi Agent Deep Reinforcement Learning
The core idea of reinforcement learning is to have an agent learning directly from exposure to an unknown environment. It is formally modeled as a Markov Decision Process (MDP) $\{S,A,r\}$ where $S$ is the state space, $A$ the action space and $r:S\times A\to\mathbb{R}$ the reward function. We define the discounted return over a trajectory of length $L$ as $R(s,a):=\sum_{t=0}^{t=L-1}\gamma^{t}r(s_{t},a_{t})$, where $0<\gamma<1$ is the discount factor. An agent interacts with the environment sampling action from its policy $a\sim\mu(s)$, and it aims at maximizing the reward, so the goal is to find the policy that maximizes the discounted return over an infinite horizon. In particular, defining the $Q$-function as:
@@ IV-B Multi Agent Deep Reinforcement Learning
we want to solve the following optimization to obtain the optimal policy $\pi^{*}$:
$Q^{*}(s,a)=Q_{\mu^{*}}(s,a)=\max_{\mu}Q_{\mu}(s,a).$
Generally, we do not have access to analytical form for neither $Q$ nor $\mu$, making the optimization extremely complex to solve. RL algorithms, and in particular deep RL (DRL) algorithms, try to solve this problem learning a policy that approximates the optimal one. DRL employs neural networks to model both the policy $\mu$ and the $Q$-function, so that the resulting algorithm is model free and can be used for continuous or discrete states and actions. We will not explore in depth here the differences between DRL algorithms, but we will provide a quick overview of the one we use.
We implemented the multi-agent deep reinforcement learning algorithm described in [21]: this is called an actor-critic algorithm, where an actor learns the policy and a critic the $Q$-function. The actor and critic are neural networks parameterized by $\theta$ and $\omega$ respectively, which are trained over specific losses. In particular the loss function for the critic is based on the Bellman’s equation:
$\mathcal{L}_{Q}=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[(Q_{\omega}(s,a)-(r+\beta Q_{\omega}(s^{\prime},\mu_{\theta}(s^{\prime}))))^{2}\right],$
@@
The intuition behind this algorithm is that the critic learns to map future rewards to specific states and actions (which is the $Q$-function), while the actor tries to maximize the $Q$-function itself (represented by the critic output). In the multi-agent settings we appply a modified version of the single agent implementation presented here, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents. This was introduced in [21] as a way of stabilizing the learning and improving the results. In particular, given the zonal nature of our environment, we share the actions between the agents of the same zone, but not between all agents. In Alg. 2 we report in pseudo-code the structure for the training of the RL agents.
We leave further details on the implementation and references to the code in Appendix E.

### v4

@@ III Theoretical Framework
In a large part of the literature, the convexity of the agents’ optimization problems is an essential assumption under which noncooperative games are studied and analyzed. On the contrary, in this paper, $\Gamma$ is a generalized Nash game with side constraints, involving trilinear objective functions and nonconvex constraints; thus, giving rise to a nonconvex game. In order to proceed, we will rely on results established in [pang_nonconvex_2011], where the authors study a class of problems with similar settings. To that purpose, we first apply the generic setting of [pang_nonconvex_2011]. Then, we present a relaxed version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version. Further, we prove existence of market equilibrium under mild assumptions, and highlight the generalized potential structure of $\Gamma$.
III-A Nonconvex Game with Side Constraints
The formulation of [pang_nonconvex_2011] considers a problem where the coupling constraints are treated as penalties with associated prices, in a modified objective function. This problem is then solved as an $(N+1)$-player game, where $N$ players optimize their variables and an external one optimizes the prices, which are shared by all players. In our formulation, the coupling constraints are Eqs. (4)-(6)-(7)-(13d)-(13f)-(13g), while the remaining part is considered as local constraints.
With the compact notation previously introduced, we can write the problem as an $(N+1)$-player game with prices
$\forall n\in\mathcal{N},\quad\max_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[J_{n}(w_{n},w_{-n})-\sum_{j=1}^{r}\rho_{j}G_{j}(w)\right],$ (15)
where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$. We can now recall the definition of Nash Equilibrium.
Definition 1 (Nash Equilibrium with Side Constraints, [pang_nonconvex_2011, Definition 1]).
A Nash equilibrium (NE) for the game (15) is a strategy-price tuple $(w^{\star},\rho^{\star})$, such that for all $n$
$w_{n}^{\star}\in\operatorname*{argmax}_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[J_{n}(w_{n},w_{-n}^{\star})-\sum_{j=1}^{r}\rho_{j}^{\star}G_{j}(w_{n},w_{-n}^{\star})\right],$
and
$0\leq\rho^{\star}\perp G(w^{\star})\leq 0.$
III-B Relaxed Generalized Nash Game
In order to extend the results from [pang_nonconvex_2011], we consider a relaxed version of our problem. We will use this relaxed version to prove the existence of an equilibrium under mild assumptions. Concretely, we add the constraints from Eqs. (4) and (7) to the objective function, which gives the function $J_{n}^{p}(w)$, the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, and the associated set of penalties $M_{n}$ for each agent. We can formulate the penalized generalized Nash game $\Gamma^{p}$ as follows:
$\max_{w_{n}\in\mathcal{W}_{n}}$ $J_{n}^{p}(w)$ (16a)
@@ III-B Relaxed Generalized Nash Game
Assume $K_{n}\geq\frac{Z}{3},\,\forall n$. Then, there exists a set of penalties $(M_{j})_{j\in\mathcal{N}}$ such that a solution of the penalized game (16) is a solution of $\Gamma$.
Proof.
We call the two Lagrangian functions associated to the problems $\mathcal{L}_{n}^{P}$ and $\mathcal{L}_{n}^{U}$, and we consider the solution for $U$: $(w_{n}^{*},\Theta_{n}^{*})\in\mbox{SOL}\{\mbox{KKT of }U\}$, where $\Theta$ is the set of Lagrangian multipliers. This solution must respect all the constraints and be a solution of $\nabla\mathcal{L}^{U}(w^{*}_{n},\Theta^{*}_{n})=0$. We now call the part of Lagrangian multipliers that are still present in $P$ $\Theta[\Sigma_{n}]$, and we consider the set of equations coming from $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma_{n}];M_{n})=0$, where $M_{n}$ is the set of penalties for agent $n$. If we evaluate these equations in $(w_{n}^{*},\Theta_{n}^{*}[\Sigma_{n}])$, we can solve them w.r.t. the penalties $M_{n}$, obtaining $\bar{M}_{n}$. Given the assumptions, solving the system $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma];\bar{M}_{n})=0$ in $(w_{n},\Theta_{n}[\Sigma])$ provides $\tilde{w}_{n}$ which is a solution also for $\nabla\mathcal{L}^{U}(w_{n},\Theta_{n})=0$. ∎
@@ III-C Existence of a Generalized Nash Equilibrium for $\Gamma$
To obtain existence result for the relaxed problem, we have to verify assumptions on the variable domains, the objective functions and the constraint sets.
@@
    $J_{n}^{p}(\cdot,w_{-n})$ is continuously differentiable in $w_{n}$, $h_{n}^{p}(\cdot)$ and $G^{p}(\cdot,w_{-n})$ are twice continuously differentiable in $w_{n}$ for all $n$;
@@
    There exists $w^{\star}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h_{n}^{p}(w_{n}^{\star})<0,\,\forall n$ and $G^{p}(w^{\star})<0$;
  • (d)
    The Hessians matrices $\nabla^{2}h_{n}^{p}(w_{n})$ for the $l_{n}$ local constraints are copositive for all $n$;
  • (e)
    The Hessians matrices $\nabla^{2}G^{p}(w)$ for the $r$ global constraints are copositive.
Theorem 1 (Existence of a solution for $\Gamma^{p}$).
Define $\mathcal{W}^{\star}\subseteq\mathcal{W}$ such that for any $w\in\mathcal{W}^{\star}$, the Hessian matrix of the objective function is positive semi-definite. Then, there exists a generalized Nash equilibrium on $\mathcal{W}^{\star}$ solution of $\Gamma$.
@@
Assumption 1 (a) holds, since $\mathcal{W}_{n}$ is defined as a polyhedral and compact set, i.e., we can express any condition of the form $x\in[x_{a},x_{b}]$ as a set of inequalities $x\leq x_{b},x\geq x_{a}$, which are defined as intersections between closed half spaces. Further, the objective function is continuous and differentiable everywhere in the domain $\mathcal{W}^{\star}$, and the constraints functions are bilinear or trilinear. Thus, they are twice continuous and differentiable over the whole domain $\mathcal{W}^{\star}$. Assumption 1 (b) is then satisfied. Assumption 1 (c) requires the existence of an interior point. Considering $w^{\star}$ that strictly satisfies Eq. (4), (5), (6) and (7), we immediately obtain that $\lambda_{n}^{z}=0,\mu_{n,k}=0,\sigma_{n}=0,\delta_{n}=0,\,\forall k\in\{1,\ldots,K_{n}\},\,\forall n\in\mathcal{N},\,\forall z\in\mathcal{Z}$. Eq. (13c) needs instead to be relaxed to a strict inequality. Finally, to check for Assumptions 1 (d)-(e), we need to compute the Hessians matrices associated to the constraints functions. Regarding the local set of constraints $h_{n}^{p}(w_{n})$, we obtain $\nabla^{2}h_{n}^{p}(w_{n})=0,\,\forall n\in\mathcal{N},\,\forall w_{n}\in\mathcal{W}_{n}$, except for Eq. (13c), for which we obtain the following cases:
Given that an interior point forces $\lambda_{n}^{z}=0,\delta_{n}=0$ and $\sigma_{n}=0$, and that the prices $\pi_{n,k}$ are positive by definition, both inequalities are always satisfied. Considering the only remaining global constraint, Eq (6), we obtain a constant positive Hessian , which directly satisfies the copositivity assumption. We then apply [pang_nonconvex_2011, Theorem 13] to the $(N+1)$-players game to obtain the existence of a Nash equilibrium $(w^{\star},\rho^{\star})$, and from [pang_nonconvex_2011, Section 2.1] we can conclude that $w^{\star}$ is a generalized Nash equilibrium for $\Gamma^{p}$. Finally, using Proposition 1, a solution of $\Gamma^{p}$ is also a solution for $\Gamma$, which concludes the proof. ∎
@@
Using [facchinei_decomposition_2011, Def. 2.1], two requirements need to be satisfied for a GNG to be a GPG. The first requirement amounts to have a feasible set that is nonempty, which in our setting implies having enough energy offer to cover the total demand. The second requirement is for a potential $P(w)$ function to exist, which in our case is satisfied by using $P(w):=\sum_{n=1}^{N}J_{n}(w_{n},w_{-n})$. ∎
From Proposition 2, we can rewrite Eqs. (14) as
@@
$\quad G(w_{n},w_{-n})\leq 0,$ (17c)
which can be solved using a standard nonlinear programming solver. The potential structure is interesting algorithmically, to guarantee the convergence of best-response dynamics.
@@ IV Algorithmic Solutions
We propose two exact approaches to compute GNEs of $\Gamma$, that we compare against MARL simulation which capture in a realistic way the interactions between the producers and the market dynamics. One of the exact approach makes use of the GPG structure of $\Gamma$; thus leading to compute a solution of (17) in an integrated way, relying on a nonlinear mathematical optimization solver (MINOS) which is based on the projected Lagrangian method [noauthor_minos_nodate]. We describe next the other exact approach.
IV-A Gauss-Seidel Best-Response
We follow the implementation of the Gauss-Seidel best-response algorithm from [facchinei_decomposition_2011]. We solve the problem in Eq. (14) for each agent sequentially (using MINOS to compute best-response for each player); stopping the whole process once the stopping criterion is met, i.e., the distance between the players’ strategy profiles at two consecutive iterations do not change more than a small constant $\varepsilon$. This is summarized in Alg. 1.
@@ IV-A Gauss-Seidel Best-Response
1:Initialize problem with feasible point $w^{0}=[w_{1}^{0},\dots,w_{N}^{0}]$, $k=0$, $\varepsilon>0$
2:while $w^{k}-w^{k-1}||>\varepsilon$ do
3:  for $n\in\{1,\dots,N\}$ do
@@
From [facchinei_decomposition_2011, Theorem 5.2] and the GPG structure of the game, the Gauss-Seidel best-response algorithm is proved to converge in finite time to a GNE of $\Gamma$.
IV-B Multi-Agent Deep Reinforcement Learning
The core idea of reinforcement learning is to have an agent learning directly from exposure to an unknown environment. It is formally modeled as a Markov Decision Process (MDP) $\{S,A,r\}$ where $S$ is the state space, $A$ the action space and $r:S\times A\to\mathbb{R}$ the reward function. We define the discounted return over a trajectory of length $L$ as $R(s,a):=\sum_{t=0}^{t=L-1}\gamma^{t}r(s_{t},a_{t})$, where $0<\gamma<1$ is the discount factor. An agent interacts with the environment sampling action from its policy $a\sim\mu(s)$, and it aims to maximize the reward, so the goal is to find the policy that maximizes the discounted return over an infinite horizon. In particular, defining the $Q$-function as:
@@ IV-B Multi-Agent Deep Reinforcement Learning
we want to solve the following optimization problem to obtain the optimal policy $\pi^{\star}$
$Q^{\star}(s,a)=Q_{\mu^{\star}}(s,a)=\max_{\mu}Q_{\mu}(s,a).$
Generally, we do not have access to analytical forms for neither $Q$ nor $\mu$, making the optimization extremely complex to solve. RL algorithms, and in particular deep RL (DRL) algorithms, try to solve this problem learning a policy that approximates the optimal one. DRL employs neural networks to model both the policy $\mu$ and the $Q$-function, so that the resulting algorithm is model free and can be used for continuous or discrete states and actions. We will not explore in depth here the differences between DRL algorithms, but we will provide a quick overview of the one we use.
We implemented the multi-agent deep reinforcement learning algorithm described in [lowe_multi-agent_2017]: this is called an actor-critic algorithm, where an actor learns the policy and a critic the $Q$-function. The actor and critic are neural networks parameterized by $\theta$ and $\omega$ respectively, which are trained over specific losses. In particular the loss function for the critic is based on the Bellman’s equation:
$\mathcal{L}_{Q}=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[(Q_{\omega}(s,a)-(r+\gamma Q_{\omega}(s^{\prime},\mu_{\theta}(s^{\prime}))))^{2}\right],$
@@
The intuition behind this algorithm is that the critic learns to map future rewards to specific states and actions (which is the $Q$-function), while the actor tries to maximize the $Q$-function itself (represented by the critic output). In the multi-agent settings we apply a modified version of the single agent implementation presented here, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents. This was introduced in [lowe_multi-agent_2017] as a way of stabilizing the learning and improving the results. In particular, given the zonal nature of our environment, we share the actions between the agents of the same zone, but not between all agents.
In our environment, the action $a$ of each producer corresponds to its bids (prices and capacities), while the state $s$ contains the previous day clearing prices, information about the current day and a random zonal perturbation. The reward is a function of the profit and clearing price. In order to run the simulation, each producer outputs their bids, which are used to clear the market solving problem (9). The results are then used to compute the reward for each producer, which in turn is used to update their strategy. In Alg. 2 we report in pseudo-code the structure for the training of the RL agents.